# Experimental Demonstration of an On-Axis Laser Ranging Interferometer for Future Gravity Missions

### v1

@@ Abstract
We experimentally demonstrate a novel interferometric architecture for next-generation gravity missions, featuring a laser ranging interferometer (LRI) that enables monoaxial transmission and reception of laser beams between two optical benches with a heterodyne frequency of 7.3 MHz. Active beam steering loops, utilizing differential wavefront sensing (DWS) signals, ensure co-alignment between the receiving (RX) beam and the transmitting (TX) beam. With spacecraft attitude jitter simulated by hexapod-driven rotations, the interferometric link achieves a pointing stability below 10 µrad/ in the frequency range between 2 mHz and 0.5 Hz, and the fluctuation of the TX beam’s polarization state results in a reduction of 0.14% in the carrier-to-noise-density ratio over a 15-hour continuous measurement. Additionally, tilt-to-length (TTL) coupling is experimentally investigated using the periodic scanning of the hexapod. Experimental results show that the on-axis LRI enables the inter-spacecraft ranging measurements with nanometer accuracy, making it a potential candidate for future GRACE-like missions.
@@ I INTRODUCTION
The Gravity Recovery and Climate Experiment (GRACE) mission [tapley2004gravity], a collaboration between the National Aeronautics and Space Administration (NASA) and the German Aerospace Center (DLR), was operational from March 2002 to October 2017. It employed twin satellites separated by approximately 200 kilometers to precisely measure inter-spacecraft distance variations via a Microwave Instrument (MWI), enabling the retrieval of Earth’s time-variable gravity field [tapley2005ggm02]. Monthly gravity field maps derived from the GRACE data were used to investigate global-scale mass transport phenomena, including ice-sheet dynamics, ocean circulation patterns, and continental hydrologic processes, which are critical for understanding the dynamics of Earth’s climate system [tapley2004grace].
@@
As the successor to the GRACE mission, the GRACE Follow-On (GRACE-FO) mission was launched in May 2018 to extend continuous monitoring of Earth’s gravity field [Kornfeld2019]. In addition to the MWI serving as the primary instrument, the mission incorporated a Laser Ranging Interferometer (LRI) as a technology demonstration. To mitigate temporal gaps between the GRACE-FO mission and its successors, future GRACE-like missions are under active development. GRACE-Continuity (GRACE-C), a joint mission between the United States and Germany, is scheduled for launch in 2028 to ensure uninterrupted data collection [bender2025short]. In parallel, the European Space Agency (ESA) is progressing with the Next-Generation Gravity Mission (NGGM), which will employ a Bender constellation to enhance spatial and temporal resolution [haagmans2020esa], with an anticipated launch in the early 2030s.
@@
The GRACE-FO LRI successfully tracked the inter-spacecraft distance with a sensitivity level of 1 nm/ at frequencies around 100 mHz, which is two orders of magnitude better than the performance of the MWI [ghobadi2020grace, abich2019orbit]. The GRACE-FO LRI utilizes optical heterodyne interferometry in a transponder configuration, representing the first demonstration in space of inter-spacecraft laser links [sheard2012intersatellite]. In the reference spacecraft, a laser is frequency-stabilized to a highly stable optical cavity. A beamsplitter diverts a small fraction of the laser to serve as a local oscillator beam, while most of the laser is transmitted to the remote spacecraft. In the transponder spacecraft, the laser is frequency-offset-locked to the incoming beam and transmitted back to the reference spacecraft [abich2019orbit, heinzel2006lisa]. The returning beam is then combined with the local beam to produce the heterodyne signal, enabling the measurement of relative phase shifts that encode the round-trip distance variation between the two spacecraft. The reference and transponder spacecraft are built identically, allowing them to switch operational roles. In addition, an active beam steering method was implemented in the LRI to mitigate the effect of the spacecraft attitude jitter [schutze2014laser]. Owing to the success of the GRACE-FO LRI, the LRI is selected to be the primary instrument in the upcoming GRACE-C and NGGM missions [landerer2024towards, nicklaus2022towards, nicklaus2020laser].
@@
The GRACE-FO LRI employs an off-axis configuration using a triple mirror assembly (TMA), which introduces a lateral offset between the receiving (RX) and transmitting (TX) beams [schutze_retroreflector_2014, ward2014design]. This design circumvents the constraint imposed by the MWI, which occupies the line of sight between the centers of mass of the two spacecraft. To suppress the tilt-to-length (TTL) coupling that arises from the spacecraft attitude jitter, the virtual vertex of the TMA is aligned with the center of mass (CoM) of the spacecraft, where the accelerometer is located. The LRI will be the main instrument in future missions. Without the accommodation constraints imposed by the MWI, it becomes feasible to investigate different topologies. Mandel et al. designed a mono-axis architecture using a retro-reflector and polarizing optics, which integrates multiple methodologies for heterodyne laser ranging while enabling beam-tracking capabilities [mandel2020architecture]. Our group has proposed an on-axis architecture for laser ranging systems, which utilizes polarization-dependent routing to maintain the anti-parallelism between the RX beam and the TX beam. This design was thoroughly validated through numerical simulations, demonstrating its potential for high-performance laser ranging [yang2022axis]. The experimental result presented here extends this framework by incorporating modifications to the proposed on-axis topology.
@@
This paper reports the experimental results of an on-axis LRI, which establishes a transponder-based laser interferometric link featuring two independent active beam steering loops. A laboratory-developed breadboard model demonstrates an on-axis interferometric link via anti-parallel RX and TX beams between two optical benches. With the angular jitters, the beam pointing stability was characterized using calibrated differential wavefront sensing (DWS) and differential power sensing (DPS) signals. The potential fluctuation of the TX beam’s polarization state was quantified, and its impact on the carrier-to-noise-density ratio (CNR) was analyzed. Additionally, the TTL coupling of the optical bench was investigated through dedicated rotational configurations based on the experimental model.
@@ II ON-AXIS LRI CONCEPT
The concept of the inter-spacecraft laser interferometry for geodesy in an on-axis configuration is shown in Fig. 1. The twin spacecraft, operating in a trailing formation, orbit at altitudes 300–500 km above Earth with an approximate separation of 200 km in a near-polar orbit. In the reference spacecraft, the laser is frequency-stabilized to an optical cavity via the Pound–Drever–Hall (PDH) technique [drever_laser_1983]. To operate in a transponder scheme, the laser in the transponder is frequency-offset-locked to the received signal from the reference spacecraft. The identical optical benches in the two spacecraft utilize polarizing optics to enable monoaxial transmission and reception of laser beams. The combination of a mirror (M1) and a quarter-wave plate (QWP) converts the perpendicular-polarized (S-pol) beam reflected by the polarizing beamsplitter (PBS) into the parallel-polarized(P-pol) light, thereby directing the majority of the laser power toward the remote spacecraft after retransmission through the PBS. In contrast to the off-axis configuration, the TX and RX reference points, which define the pivot points to achieve the minimal TTL coupling for the TX beam and the RX beam [Muller2017PhD], are co-located in the aperture of the optical bench. The fast steering mirror (FSM), the aperture, and the quadrant photoreceivers (QPRs) are arranged to establish optical conjugate planes using imaging systems, making the path length insensitive to the spacecraft’s angular jitter. The DWS signals related to the spacecraft’s pointing angle are fed back to the FSM or the attitude and orbit control system to maintain link strength and mitigate TTL coupling.
@@
The off-axis configuration, characterized by a simplified optical bench and a high-precision TMA, does not utilize polarizing components. As evidenced by the GRACE-FO optical bench, this design incorporates a minimal number of critical components, reducing the stability requirements for the optical bench. However, its inherent racetrack configuration complicates telescope integration, thereby limiting the maximum measurable distance between spacecraft. In contrast, the on-axis topology leverages polarizing optics with a shared reference point on the optical bench, utilizing imaging systems and mirrors to achieve anti-parallel transmission and reception. This design minimizes the TTL coupling caused by angular jitter while eliminating the requirement for a TMA, a critical component in off-axis systems. The on-axis topology enables the utilization of a common telescope to enhance the received beam power and reduce the divergence of the TX beam, which is particularly beneficial for shot-noise-limited interferometers like the Laser Interferometer Space Antenna (LISA) mission [LISA]. For example, in Fig. 1, the telescope can be placed between the aperture and the folding mirrors. Moreover, the on-axis topology facilitates beam transmission and reception using a single telescope and a single baffle, resulting in a more compact and space-efficient system.
@@ III.1 Experimental setup
Fig. 3 shows the experimental setup, which consists of a reference bench, a transponder bench, and a phasemeter. As a modification of the concept we proposed in [yang2022axis], the reference bench is installed in a hexapod (Newport, HXP100-MECA), which provides six-degrees-of-freedom motions to simulate spacecraft attitude jitter in the experiment. In this experiment, the definitions of the RX beam and the TX beam are referenced with the reference bench. The transponder bench is fixed on an optical table and used to produce the RX beam, simulating the far-field beam from a remote spacecraft. The phasemeter tracks the phase of beat-note signals, which is then used to implement the active beam steering loops. The transponder scheme, validated in the GRACE-FO mission [abich2019orbit, kornfeld2019grace, sheard2012intersatellite], is implemented in this experiment, where the transponder laser is frequency-offset locked with the laser from the reference bench. As enlarged excerpts from the reference bench in Fig. 3, Fig. 4 illustrates the polarization states of beams and highlights the principle of the beam steering loop used in this experiment.
@@
The RX beam is reflected off a mirror M5 and enters the system via a clipping at the aperture RX(TX)RP. By orienting the fast axis at 45∘, QWP2 transforms the P-pol LO beam and the S-pol RX beam into right-handed and left-handed circularly polarized beams, thereby enabling the heterodyne interference between the LO beam and the RX beam at PBS2. The beat-note beams in orthogonal polarization states are split by PBS2 and then detected by two InGaAs four-quadrant photoreceivers, QPR1 and QPR2, respectively. A half-wave plate (HWP1) is inserted after PBS1, with its fast axis oriented at 45∘, which converts both the TX and RX beams to S polarization. Unlike the 45∘ linear polarization of the TX beam used in [yang2022axis], the configuration of pure S polarization of the TX beam prevents potential phase retardance caused by subsequent reflections, thereby eliminating the need for unpolarized coatings on subsequent mirrors. All waveplates, including QWP1, QWP2, and HWP1, are true zero-order waveplates, minimizing the TTL coupling due to phase retardance [gu2017study].
@@
On the transponder bench, a laser beam from a fiber collimator is split into two parts by a beam splitter BS3. The smaller portion is guided to the single-element photoreceiver (SEPR) as a reference beam, enabling heterodyne detection with the TX beam from the reference bench. The phasemeter, employing all-digital phase-locked loops (ADPLLs), measures the beat-note frequency of the SEPR signal [gerberding2013phasemeter, shaddock2006overview]. This frequency is fed back to a fast piezoelectric transducer (PZT) actuator and a slow temperature controller to regulate the laser frequency of the transponder laser [shaddock2006overview, schwarze2018phase], achieving a 7.3 MHz frequency offset lock. The open-loop transfer function of the laser frequency lock was measured using a frequency response analyzer (Moku:Lab, Liquid Instruments). As shown in Fig. 5, the laser frequency lock loop exhibits a unity-gain bandwidth of 32.36 kHz and a phase margin of 44.67∘.
@@
The phasemeter, originally developed for the Laser Interferometer Space Antenna (LISA), is capable of tracking nanometer translations and nanoradian-scale tilts from the heterodyne signals of two laser beams. The phase readout of the heterodyne signals is based on ADPLLs [gerberding2013phasemeter, shaddock2006overview], which are implemented in a field-programmable gate array (FPGA) platform. By combining the phase information from four individual channels, each connecting the signals from the segments of a QPR, the phasemeter calculates and processes the horizontal and vertical DWS signals [heinzel2020tracking]. These DWS signals characterize the relative wavefront tilt of the two laser beams and are employed in the beam steering in this experiment.
@@ III.2 DWS and DPS
The DWS signals acquired from QPR1 and QPR2 encode the relative wavefront tilt between the RX beam and the TX beam. The horizontal and vertical DWS signals are linearly proportional to the yaw and pitch misalignment over a small angular range [wanner2012methods, heinzel2020tracking, morrison1994automatic]. With misalignment angles (yaw) and (pitch), the DWS signals are derived via the conversion matrix ,
where and denote the horizontal and vertical DWS signals, respectively. The diagonal elements and represent direct coupling factors between the misalignment angles and the corresponding DWS signals. The off-diagonal elements and describe cross-coupling effects, wherein yaw misalignment induces vertical DWS signals and pitch misalignment results in horizontal DWS signals. Intrinsic DWS offsets arising from inter-segment variations in readout electronics, cable lengths, and QPR misalignment were observed in QPR-based measurements. These offsets were quantified by a modulated optical signal, which was artificially generated by amplitude-modulating the LO beam while blocking the RX beam. Throughout this paper, the measured DWS signals refer to the signals preprocessed by offset subtraction.
@@
To assess the stability of the beam steering loops, the open-loop transfer functions were characterized using a frequency response analyzer (Moku:Lab, Liquid Instruments). As shown in Fig. 8, the control loop based on the horizontal DWS signal exhibits a unity-gain bandwidth of 141.75 Hz and a phase margin of 55.01∘, whereas the control loop based on the vertical DWS signal demonstrates a unity-gain bandwidth of 108.72 Hz and a phase margin of 45.90∘. The bandwidths of both loops are much higher than the frequency of spacecraft attitude jitter, which is below 1 Hz in general [daniel_intersatellite_2015]. Moreover, the phase margins of both loops exceed the typical threshold of 45∘, indicating that both beam steering loops can operate with adequate robustness.

### v2

@@ Abstract
We experimentally demonstrate a novel interferometric architecture for next-generation gravity missions, featuring a laser ranging interferometer (LRI) that enables monoaxial transmission and reception of laser beams between two optical benches with a heterodyne frequency of 7.3 MHz. Active beam steering loops, utilizing differential wavefront sensing (DWS) signals, ensure co-alignment between the receiving (RX) beam and the transmitting (TX) beam. With spacecraft attitude jitter simulated by hexapod-driven rotations, the interferometric link achieves a pointing stability below 10 µrad/ in the frequency range between 0.2 mHz and 0.5 Hz, and the fluctuation of the TX beam’s polarization state results in a reduction of 0.14% in the carrier-to-noise-density ratio over a 15-hour continuous measurement. Additionally, tilt-to-length (TTL) coupling is experimentally investigated using the periodic scanning of the hexapod. Experimental results show that the on-axis LRI enables the inter-spacecraft ranging measurements with nanometer accuracy, making it a potential candidate for future GRACE-like missions.
@@ I INTRODUCTION
The Gravity Recovery and Climate Experiment (GRACE) mission [35], a collaboration between the National Aeronautics and Space Administration (NASA) and the German Aerospace Center (DLR), was operational from March 2002 to October 2017. It employed twin satellites separated by approximately 200 kilometers to precisely measure inter-spacecraft distance variations via a Microwave Instrument (MWI), enabling the retrieval of Earth’s time-variable gravity field [33]. Monthly gravity field maps derived from the GRACE data were used to investigate global-scale mass transport phenomena, including ice-sheet dynamics, ocean circulation patterns, and continental hydrologic processes, which are critical for understanding the dynamics of Earth’s climate system [34].
@@
As the successor to the GRACE mission, the GRACE Follow-On (GRACE-FO) mission was launched in May 2018 to extend continuous monitoring of Earth’s gravity field [16]. In addition to the MWI serving as the primary instrument, the mission incorporated a Laser Ranging Interferometer (LRI) as a technology demonstration. To mitigate temporal gaps between the GRACE-FO mission and its successors, future GRACE-like missions are under active development. GRACE-Continuity (GRACE-C), a joint mission between the United States and Germany, is scheduled for launch in 2028 to ensure uninterrupted data collection [4]. In parallel, the European Space Agency (ESA) is progressing with the Next-Generation Gravity Mission (NGGM), which will employ a Bender constellation to enhance spatial and temporal resolution [11], with an anticipated launch in the early 2030s.
@@
The GRACE-FO LRI successfully tracked the inter-spacecraft distance with a sensitivity level of 1 nm/ at frequencies around 100 mHz, which is two orders of magnitude better than the performance of the MWI [8, 1]. The GRACE-FO LRI utilizes optical heterodyne interferometry in a transponder configuration, representing the first demonstration in space of inter-spacecraft laser links [32]. In the reference spacecraft, a laser is frequency-stabilized to a highly stable optical cavity. A beamsplitter diverts a small fraction of the laser to serve as a local oscillator beam, while most of the laser is transmitted to the remote spacecraft. In the transponder spacecraft, the laser is frequency-offset-locked to the incoming beam and transmitted back to the reference spacecraft [1, 13]. The returning beam is then combined with the local beam to produce the heterodyne signal, enabling the measurement of relative phase shifts that encode the round-trip distance variation between the two spacecraft. The reference and transponder spacecraft are built identically, allowing them to switch operational roles. In addition, an active beam steering method was implemented in the LRI to mitigate the effect of the spacecraft attitude jitter [29]. Owing to the success of the GRACE-FO LRI, the LRI is selected to be the primary instrument in the upcoming GRACE-C and NGGM missions [17, 27, 25].
@@
The GRACE-FO LRI employs an off-axis configuration using a triple mirror assembly (TMA), which introduces a lateral offset between the receiving (RX) and transmitting (TX) beams [28, 39]. This design circumvents the constraint imposed by the MWI, which occupies the line of sight between the centers of mass of the two spacecraft. To suppress the tilt-to-length (TTL) coupling that arises from the spacecraft attitude jitter, the virtual vertex of the TMA is aligned with the center of mass (CoM) of the spacecraft, where the accelerometer is located. The LRI will be the main instrument in future missions. Without the accommodation constraints imposed by the MWI, it becomes feasible to investigate different topologies. Mandel et al. designed a mono-axis architecture using a retro-reflector and polarizing optics, which integrates multiple methodologies for heterodyne laser ranging while enabling beam-tracking capabilities [21]. Our group has proposed an on-axis architecture for laser ranging systems, which utilizes polarization-dependent routing to maintain the anti-parallelism between the RX beam and the TX beam. This design was thoroughly validated through numerical simulations, demonstrating its potential for high-performance laser ranging [43]. The experimental result presented here extends this framework by incorporating modifications to the proposed on-axis topology.
@@
This paper reports the experimental results of an on-axis LRI, which establishes a transponder-based laser interferometric link featuring two independent active beam steering loops. A laboratory-developed breadboard model demonstrates an on-axis interferometric link via anti-parallel RX and TX beams between two optical benches. With the angular jitters, the beam pointing stability was characterized using calibrated differential wavefront sensing (DWS) and differential power sensing (DPS) signals. The potential fluctuation of the TX beam’s polarization state was quantified, and its impact on the carrier-to-noise-density ratio (), defined as the ratio of carrier power to noise power spectral density, was analyzed. Additionally, the TTL coupling of the optical bench was investigated through dedicated rotational configurations based on the experimental model.
Beyond the GRACE-like missions, the concepts and technologies demonstrated in this work could have broader application potential. In the field of gravitational wave detection, the Laser Interferometer Space Antenna (LISA) mission will establish three laser interferometer links between three spacecraft arranged in a triangular array [2]. Related space-based initiatives include the Chinese Taiji [20] and Tianqin [19] projects, as well as the Japanese DECi-hertz Interferometer Gravitational wave Observatory (DECIGO) [14]. Additionally, the precision beam steering and DWS technologies developed here are applicable to laser communication systems, where maintaining precise alignment is critical for minimizing link loss and maximizing data throughput. Moreover, the space interferometric technology explored here could facilitate fundamental experiments in space quantum interference [37, 42, 18].
@@ II ON-AXIS LRI CONCEPT
The concept of the inter-spacecraft laser interferometry for geodesy in an on-axis configuration is shown in Fig. 1. The twin spacecraft, operating in a trailing formation, orbit at altitudes 300–500 km above Earth with an approximate separation of 200 km in a near-polar orbit. In the reference spacecraft, the laser is frequency-stabilized to an optical cavity via the Pound–Drever–Hall (PDH) technique [6]. To operate in a transponder scheme, the laser in the transponder is frequency-offset-locked to the received signal from the reference spacecraft. The identical optical benches in the two spacecraft utilize polarizing optics to enable monoaxial transmission and reception of laser beams. The combination of a mirror (M1) and a quarter-wave plate (QWP) converts the perpendicular-polarized (S-pol) beam reflected by the polarizing beamsplitter (PBS) into the parallel-polarized(P-pol) light, thereby directing the majority of the laser power toward the remote spacecraft after retransmission through the PBS. In contrast to the off-axis configuration, the TX and RX reference points, which define the pivot points to achieve the minimal TTL coupling for the TX beam and the RX beam [23], are co-located in the aperture of the optical bench. The fast steering mirror (FSM), the aperture, and the quadrant photoreceivers (QPRs) are arranged to establish optical conjugate planes using imaging systems, making the path length insensitive to the spacecraft’s angular jitter. The DWS signals related to the spacecraft’s pointing angle are fed back to the FSM or the attitude and orbit control system to maintain link strength and mitigate TTL coupling.
@@
The off-axis configuration, characterized by a simplified optical bench and a high-precision TMA, does not utilize polarizing components. As evidenced by the GRACE-FO optical bench, this design incorporates a minimal number of critical components, reducing the stability requirements for the optical bench. However, its inherent racetrack configuration complicates telescope integration, thereby limiting the maximum measurable distance between spacecraft. In contrast, the on-axis topology leverages polarizing optics with a shared reference point on the optical bench, utilizing imaging systems and mirrors to achieve anti-parallel transmission and reception. This design minimizes the TTL coupling caused by angular jitter while eliminating the requirement for a TMA, a critical component in off-axis systems. The on-axis topology enables the utilization of a common telescope to enhance the received beam power and reduce the divergence of the TX beam, which is particularly beneficial for shot-noise-limited interferometers like the LISA mission [2]. For example, in Fig. 1, the telescope can be placed between the aperture and the folding mirrors.
As discussed in [43], the on-axis optical bench, using the same laser configuration and incorporating telescope imaging systems, shows an improvement in the , a key factor limiting the phase readout sensitivity, compared to the GRACE-FO optical bench. Moreover, the on-axis topology enables a more streamlined design, allowing for beam transmission and reception through a single telescope and baffle, which yields a more compact and space-efficient system. In contrast, off‑axis configurations such as the GRACE‑FO LRI use a retro‑reflector mounted separately from the optical bench, which increases assembly complexity and thermal‑control requirements. Because of its telescope adaptability, the on-axis topology offers distinct advantages for future potential space‑laser‑interferometry missions that require a longer interferometric arm.
@@ III.1 Experimental setup
Fig. 3 shows the experimental setup, which consists of a reference bench, a transponder bench, and a phasemeter. As a modification of the concept we proposed in [43], the reference bench is installed in a hexapod (Newport, HXP100-MECA), which provides six-degrees-of-freedom motions to simulate spacecraft attitude jitter in the experiment. In this experiment, the definitions of the RX beam and the TX beam are referenced with the reference bench. The transponder bench is fixed on an optical table and used to produce the RX beam, simulating the far-field beam from a remote spacecraft. The phasemeter tracks the phase of beat-note signals, which is then used to implement the active beam steering loops. The transponder scheme, validated in the GRACE-FO mission [1, 15, 32], is implemented in this experiment, where the transponder laser is frequency-offset locked with the laser from the reference bench. As enlarged excerpts from the reference bench in Fig. 3, Fig. 4 illustrates the polarization states of beams and highlights the principle of the beam steering loop used in this experiment.
@@
The RX beam is reflected off a mirror M5 and enters the system via a clipping at the aperture RX(TX)RP. By orienting the fast axis at 45∘, QWP2 transforms the P-pol LO beam and the S-pol RX beam into right-handed and left-handed circularly polarized beams, thereby enabling the heterodyne interference between the LO beam and the RX beam at PBS2. The beat-note beams in orthogonal polarization states are split by PBS2 and then detected by two InGaAs four-quadrant photoreceivers, QPR1 and QPR2, respectively. A half-wave plate (HWP1) is inserted after PBS1, with its fast axis oriented at 45∘, which converts both the TX and RX beams to S polarization. Unlike the 45∘ linear polarization of the TX beam used in [43], the configuration of pure S polarization of the TX beam prevents potential phase retardance caused by subsequent reflections, thereby eliminating the need for unpolarized coatings on subsequent mirrors. All waveplates, including QWP1, QWP2, and HWP1, are true zero-order waveplates, minimizing the TTL coupling due to phase retardance [10].
@@
On the transponder bench, a laser beam from a fiber collimator is split into two parts by a beam splitter BS3. The smaller portion is guided to the single-element photoreceiver (SEPR) as a reference beam, enabling heterodyne detection with the TX beam from the reference bench. The phasemeter, employing all-digital phase-locked loops (ADPLLs), measures the beat-note frequency of the SEPR signal [7, 31]. This frequency is fed back to a fast piezoelectric transducer (PZT) actuator and a slow temperature controller to regulate the laser frequency of the transponder laser [31, 30], achieving a 7.3 MHz frequency offset lock. The open-loop transfer function of the laser frequency lock was measured using a frequency response analyzer (Moku:Lab, Liquid Instruments). As shown in Fig. 5, the laser frequency lock loop exhibits a unity-gain bandwidth of 32.36 kHz and a phase margin of 44.67∘.
@@
Although a beam with lower wavefront error and intensity drop could be achieved with increasing complexity of the system, this was not considered necessary within the scope of this investigation. Even with a 40% intensity drop in the RX beam, the remained sufficient to ensure correct phase tracking by the phasemeter during this experiment. Throughout the experiment, only rotations of the reference bench around the CoM occurred, which induced a beam tilt at the RX (TX) RP. The corresponding longitudinal path length change resulting from this tilt constitutes the TTL coupling discussed in Section IV.4. While wavefront fluctuations in the RX beam introduced an initial offset in the DWS signal, this offset was calibrated using the method described in Section III.2.
@@
The phasemeter, originally developed for the LISA mission, is capable of tracking nanometer translations and nanoradian-scale tilts from the heterodyne signals of two laser beams. The phase readout of the heterodyne signals is based on ADPLLs [7, 31], which are implemented in a field-programmable gate array (FPGA) platform. By combining the phase information from four individual channels, each connecting the signals from the segments of a QPR, the phasemeter calculates and processes the horizontal and vertical DWS signals [12]. These DWS signals characterize the relative wavefront tilt of the two laser beams and are employed in the beam steering in this experiment.
@@ III.2 DWS and DPS
The DWS signals acquired from QPR1 and QPR2 encode the relative wavefront tilt between the RX beam and the TX beam. The horizontal and vertical DWS signals are linearly proportional to the yaw and pitch misalignment over a small angular range [38, 12, 22]. With misalignment angles (yaw) and (pitch), the DWS signals are derived via the conversion matrix ,
where and denote the horizontal and vertical DWS signals, respectively. The diagonal elements and represent direct coupling factors between the misalignment angles and the corresponding DWS signals. The off-diagonal elements and describe cross-coupling effects, wherein yaw misalignment induces vertical DWS signals and pitch misalignment results in horizontal DWS signals. Intrinsic DWS offsets arising from inter-segment variations in wavefront errors, readout electronics, cable lengths, and QPR misalignment were observed in QPR-based measurements. These offsets were quantified by a modulated optical signal, which was artificially generated by amplitude-modulating the RX beam while blocking the LO beam. Throughout this paper, the measured DWS signals refer to the signals preprocessed by offset subtraction.
@@
To assess the stability of the beam steering loops, the open-loop transfer functions were characterized using a frequency response analyzer (Moku:Lab, Liquid Instruments). As shown in Fig. 8, the control loop based on the horizontal DWS signal exhibits a unity-gain bandwidth of 141.75 Hz and a phase margin of 55.01∘, whereas the control loop based on the vertical DWS signal demonstrates a unity-gain bandwidth of 108.72 Hz and a phase margin of 45.90∘. The bandwidths of both loops are much higher than the frequency of spacecraft attitude jitter, which is below 1 Hz in general [5]. Moreover, the phase margins of both loops exceed the typical threshold of 45∘, indicating that both beam steering loops can operate with adequate robustness.